A body moves in a straight line and its displacement, s cm , from a fixed point P after a time , t, is given by the formula After t = 2 seconds, find its:

a) Displacement from P.

b) Velocity

c) Acceleration


1
Expert's answer
2021-09-09T13:46:02-0400

Given the displacement formula is S=t3-5t2+3t+4

When t=2s

a)

S=t3-5t2+3t+4

=23-5(22)+3(2)+4

=-2cm

b)

v="\frac{dS}{dt}"

=3t2-10t+3

given t=2s

v=3(22)-10(2)+3

=-5cm/s

c)

a="\frac{dv}{dt}"

=6t-10

t=2s

a=6(2)-10

=2m/s2